# Primal-Dual iLQR

### 13 Mar 2024 v3

@@ VI Algorithm Overview
            .gradient().eval(x[:,i], u[:,i])
        A_i = dyn_grad.transpose()[:,:n]
        B_i = dyn_grad.transpose()[:,n:]
@@
def lqr_dual_solve(lqr, dx, du):
    N = lqr.num_stages()
    dv[:,N] = Qxx * dx[:,N] + lx
    for i in reversed(range(N)):
        Qxx = lqr.Q(i).xx()
        Qxu = lqr.Q(i).xu()
        At = lqr.A(i).transpose()
        lx = lqr.l(i).x()
        dv[:,i] = Qxx * dx[:,i] + Qxu * du[:,i]
            + At * dv[:,i+1] + lx
    return dv
@@
    d = problem.total_constraints().eval(x, u)
    return g + v.transpose() * d
@@
        slope += lqr.q(i).transpose() * dxu
        slope += dv[:, i].transpose() * lqr.d(i)
    slope += lqr.q(N).x().transpose() * dx[:, N]
    slope += dv[:, N].transpose() * lqr.d(N)
@@
    m_o = eval_merit(problem, x, u, v, rho)
@@
        if m_n < m_o + kArmijo * alpha * slope:
@@
        dx, du = lqr.Solve()
        dv = lqr_dual_solve(lqr, dx, du)
@@
VII Benchmarks
@@ VIII Conclusion
This framework can easily be extended to also handle constrained problems, ideally using an augmented Lagrangian [5] or interior point [12] method; using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.

### 09 Apr 2024 v4

@@ XI Algorithm Overview
            .gradient()
            .eval(x[:,i], u[:,i])
        A_i = dyn_grad.T[:,:n]
        B_i = dyn_grad.T[:,n:]
@@
def lqr_dual_solve(P, p, dx):
    Np1 = dx.shape[1]
    return [P[i] * dx[:, i] + p[i]
        for i in range(Np1)]
@@
    d = problem.total_constraints()
        .eval(x, u)
    return g + v.T * d
@@
        slope += lqr.q(i).T * dxu
        slope += dv[:, i].T * lqr.d(i)
    slope += lqr.q(N).x().T * dx[:, N]
    slope += dv[:, N].T * lqr.d(N)
@@
    m_o = eval_merit(
        problem, x, u, v, rho)
@@
        if m_n < m_o +
            kArmijo * alpha * slope:
@@
        dx, du, P, p = lqr.Solve()
        dv = lqr_dual_solve(P, p, dx)
@@
XII Benchmarks
Figure 1: The solution of the quad-pendulum problem.
@@ XIII Conclusion
This framework can easily be extended to also handle constrained problems. Ideally, an augmented Lagrangian [5] method would be used to handle general equality constraints and an interior point [12] method would be used to handle general inequality constraints. It would also be possible to handle both general equality and inequality constraints with an augmented Lagrangian method. Using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.